# Científicas (CSIC) and Universitat de València (UV), 46980 Valencia, Spain

### v1

@@ 4.2 Weakly interacting massive particles
Figure 11: FCC-ee 95% CL exclusion for scalar WIMPs without hypercharge.
Figure 12: FCC-ee 95% CL exclusion for fermionic WIMPs without hypercharge.
@@
The matching conditions for an $SU(2)_{L}$ $n$-tuplet WIMP with hypercharge $Y$ to the SILH basis operators as normalised in Eqs. 18, 19 and 26 are
$\mathcal{S}_{2B}$ $=\frac{g_{1}^{2}}{16\pi^{2}}\frac{nY^{2}}{30M_{\rm WIMP}^{2}}N_{2},$ $\mathcal{S}_{2W}$ $=\frac{g_{2}^{2}}{16\pi^{2}}\frac{n(n^{2}-1)}{360M_{\rm WIMP}^{2}}N_{2},$ $\mathcal{S}_{3W}=\frac{g_{3}^{3}}{16\pi^{2}}\frac{n(n^{2}-1)}{2160M_{\rm WIMP}^{2}}N_{3}\,,$ (34)
where the counting factors are $N_{2}=1/2,1,4,8$ and $N_{3}=1/2,1,-1,-2$ for real scalars, complex scalars, Majorana fermions, and Dirac fermions, respectively. As discussed in the above EFT analysis, these operators can be constrained by both off-pole and on-pole data where we expect similar sensitivity and an improved combined bound. This is shown in Figs. 11 and  12 for real and complex scalar WIMPs and Majorana and Dirac fermion WIMPs, with zero hypercharge. The blue and orange bars denote the projected 95 % CL sensitivity of off-pole and on-pole data to the mass of the WIMP in GeV for various $n$-plets ranging from $n=2$ to 7. The green bar is the combined sensitivity reach, with the black hatched bars indicating the HL-LHC bounds from the study in DiLuzio:2018jwd . We see that in all cases FCC-ee can significantly extend the HL-LHC exclusions.
@@ 4.3 Custodial weak quadruplet
Figure 13: Left: FCC-ee sensitivity at 68% CL to the custodial weak quadruplet model. For comparison, the dashed blue line shows the off-pole sensitivity considering only the tree-level matching to $|H|^{6}$. Right: The black line in the $C_{H\Box}$ vs $C_{H}$ plane corresponds to the CQ model.
@@
where $\lambda_{\rm SM}$ is the SM Higgs quartic. The operator $\partial_{\mu}|H|^{2}\partial^{\mu}|H|^{2}$ is equivalent to $-Q_{H\Box}$ via integration by parts. Furthermore, the terms containing $D^{2}H$ can be reduced using the Higgs equation of motion to another $|H|^{6}$ term with coefficient $2\lambda_{\rm SM}\lambda_{4}^{2}/(3\pi^{2}M_{4}^{2})$ plus dimension-6 Yukawa corrections of the form $\sum_{\psi}Y_{\rm\psi}^{ij}|H|^{2}\bar{\psi}_{L}^{i}H\psi_{R}^{j}$. These terms correct Higgs decays, but we drop them in what follows since they are always suppressed by small SM Yukawa couplings $Y_{\rm\psi}$. Similarly, the total loop correction to $|H|^{6}$ proportional to $\lambda_{\rm SM}$ is of size $10^{-2}$ relative to the tree contribution and can be safely neglected. Thus for phenomenological purposes we may work in the limit where we keep matching contributions only in $\lambda_{4},g_{1}$, and $g_{2}$. We note that the $\lambda_{4}$ terms of our result agree with those of Durieux:2022hbu , while the contributions to $\mathcal{O}_{2B,2W,3W}$ can be easily derived from Eq. 34.
In Fig. 13 (left), we show the sensitivity of FCC-ee to the custodial weak quadruplet model on and off the $Z/W$ pole. We compare this with the off-pole sensitivity obtained only by considering the tree-level modification of the Higgs self-coupling (dashed blue line). We see that full off-pole result shows a significant reduction in sensitivity. This is because while the contribution to $\mathcal{C}_{H\Box}$ is loop suppressed and $\mathcal{C}_{H}$ is generated at tree level, $\mathcal{C}_{H\Box}$ enters $\sigma(ZH)$ at LO while $\mathcal{C}_{H}$ enters at NLO, with a relative minus sign. The CQ model predicts a fixed ratio of $\mathcal{C}_{H}/\mathcal{C}_{H\Box}=+8\pi^{2}/3$, so the two contributions partially cancel in the $ZH$ cross section. This can be seen in Fig. 13 (right), where the CQ model accidentally approximately aligns with the flattest off-pole direction. On the other hand, the $Z$-pole provides a strong bound in this direction, dominantly via the NLO contribution of $\mathcal{C}_{H\Box}$ to $\mathcal{C}_{HD}$. Finally, the $\lambda_{4}$-independent lower bound of $M_{4}\gtrsim 800$ GeV comes from the $\hat{W}$ and $\hat{Y}$ parameters, which scale as $g_{\rm EW}^{2}/M_{4}^{2}$.
@@
For $\lambda_{4}=1$ and $M_{4}=1$ TeV, this yields a contribution to the $T$ parameter of $\hat{T}=-v^{2}\mathcal{C}_{HD}(m_{Z})/2\approx 10^{-5}$, which is at the detectable level according to Eq. 22. However, $Q_{HD}$ is not the only operator activated (see e.g. Eq. 10), and using Eq. 33 we find $M_{4}/\lambda_{4}\gtrsim 1.5$ TeV, in agreement with the on-pole in Fig. 13.
@@ 5 Discussion and conclusions
We have emphasised in this work how high-precision accuracy at Tera-$Z$ complements and can even beat high-energy measurements. When they are sensitive to the same new physics modifications entering at different orders or with energy-growing effects, the increased statistics compensate for higher loop suppression on the $Z$ pole or for energy enhancements off-pole. It is a simple counting exercise to see that $Z$-pole sensitivity at NLO is of the same order of magnitude as off-pole sensitivity at LO. Tera-$Z$ is therefore expected to provide competitive NLO sensitivity to off-pole measurements at LO for any Wilson coefficient that they both have access to. This “accuracy complements energy” principle also applies at NNLO on-pole vs. NLO off-pole, as well as if a Wilson coefficient enters at LO both on-pole and off-pole but is energy-enhanced in the latter.
@@
We studied new physics contributions to the Higgs 2-point and 3-point functions, which modify the Higgs self-energy, self-coupling, and Higgs-vector-vector interactions, and to the gauge boson 2-point and 3-point functions which enter in the oblique parameters and anomalous triple-gauge boson couplings. The Higgs self-energy and vector-vector interactions contribute at leading order to $e^{+}e^{-}\to ZH$, while the Higgs self-coupling enters this process at NLO. On the other hand, they all affect the $Z$ pole oblique parameters at NLO and NNLO respectively, leading as expected to competitive on-pole projected sensitivities at FCC-ee and complementary bounds when combined with off-pole data. We find a similar story for the anomalous triple-gauge couplings that enter at leading order in $WW$ and at NLO at the $Z$ pole. Finally, the $Z$ pole oblique parameters $\hat{W}$ and $\hat{Y}$ can be constrained at leading order both on- and off-pole, with the former being statistics-enhanced and the latter energy-enhanced, such that they also lead to an overall similar and complementary projected bounds.
@@
Since so many Wilson coefficients enter the oblique parameters at higher orders, global SMEFT fits beyond leading order will find large correlations between seemingly unrelated off-pole Wilson coefficients, and cancellations between them when marginalising will give the appearance of weaker bounds. Hence it is crucial to emphasise that with Tera-$Z$ it will be especially important to match to explicit UV models to extract useful information, as the dynamical information from the model will correlate operators and pick out specific directions in the SMEFT Wilson coefficient space that typically result in stronger bounds instead. The fundamental reason is because the volume of the region that can be constrained in Wilson coefficient space is reduced by combining datasets beyond leading order. We find that this is indeed the case for the $Z_{2}$-symmetric real singlet scalar, scalar and fermion WIMPs, and a custodial weak quadruplet scalar. These explicit UV models all benefit significantly from the inclusion of on-pole data expanding coverage of their projected parameter space at FCC-ee.
@@ Appendix A SMEFT theory expression for $\sigma(e^{+}e^{-}\rightarrow W^{+}W^{-}$)
[…]
@@
$E_{\rm{CM}}$ [GeV] SM [$\mathrm{pb}$] $\mathcal{C}_{W}$ $\mathcal{C}_{HWB}$ $\mathcal{C}_{HD}$ $[\mathcal{C}_{He}]_{11}$ $[\mathcal{C}_{Hl}^{(1)}]_{11}$ $[\mathcal{C}_{Hl}^{(3)}]_{11}$ $[\mathcal{C}_{Hl}^{(3)}]_{22}$ $[\mathcal{C}_{ll}]_{1221}$
@@
[…]
As described in Sec. 3.2, it is sometimes more convenient to work in the SILH basis. Both share a large number of coefficients, but motivated by Higgs and gauge coupling modifications, as well as the situation described in Stefanek:2024kds , we focus on the subset of coefficients introduced in Eqs. 16-19 and 24 -26 and App. B.2. With these assumptions, the only operators affecting the diboson production cross section:
@@
[…]
$E_{\rm{CM}}$ [GeV] SM [$\mathrm{pb}$] $\mathcal{S}_{3W}$ $\mathcal{S}_{HB}$ $\mathcal{S}_{HW}$ $\mathcal{S}_{2B}$ $\mathcal{S}_{2W}$ $\mathcal{S}_{B+W}$ $\mathcal{S}_{B-W}$ $\mathcal{S}_{T}$
@@ Appendix B Single operator sensitivities
[…]
@@
with $\mathcal{C}_{j}$ evaluated at the $2\sigma$ limit. This can be interpreted as the observable that most strongly constrains the coefficient. For the combined sensitivity, since on- and off-pole are not correlated, the strongest bound gives the lagrest pull as well.

### v2

@@ 4.2 Weakly interacting massive particles
Figure 11: FCC-$e$ 95% CL exclusion for scalar WIMPs without hypercharge.
Figure 12: FCC-$e$ 95% CL exclusion for fermionic WIMPs without hypercharge.
@@
The matching conditions for an $SU(2)_{L}$ $n$-tuplet WIMP with hypercharge $Y$ to the SILH basis operators as normalised in Eqs. 19, 20 and 27 are
$\mathcal{S}_{2B}$ $=\frac{g_{1}^{2}}{16\pi^{2}}\frac{nY^{2}}{30M_{\rm WIMP}^{2}}N_{2},$ $\mathcal{S}_{2W}$ $=\frac{g_{2}^{2}}{16\pi^{2}}\frac{n(n^{2}-1)}{360M_{\rm WIMP}^{2}}N_{2},$ $\mathcal{S}_{3W}=\frac{g_{2}^{3}}{16\pi^{2}}\frac{n(n^{2}-1)}{2160M_{\rm WIMP}^{2}}N_{3}\,$, (35)
where the counting factors are $N_{2}=1/2,1,4,8$ and $N_{3}=1/2,1,-1,-2$ for real scalars, complex scalars, Majorana fermions, and Dirac fermions, respectively. As discussed in the above EFT analysis, these operators can be constrained by both above-pole and on-pole data where we expect similar sensitivity and an improved combined bound. This is shown in Figs. 11 and  12 for real and complex scalar WIMPs and Majorana and Dirac fermion WIMPs, with zero hypercharge. In all cases, we run down from $\Lambda=$1\text{\,}\mathrm{TeV}$\$ to the IR scale of the corresponding observable. The blue and orange bars denote the projected 95 % CL sensitivity of above-pole and on-pole data to the mass of the WIMP in GeV for various $n$-plets ranging from $n=2$ to 7. The green bar is the combined sensitivity reach, with the black hatched bars indicating the HL-LHC bounds from the study in DiLuzio:2018jwd . We see that in all cases FCC-ee can significantly extend the HL-LHC exclusions.
@@ 4.3 Custodial weak quadruplet
Figure 13: Left: FCC-$e$ sensitivity at 68% CL to the custodial weak quadruplet model. For comparison, the dashed blue line shows the above-pole sensitivity considering only the tree-level matching to $|H|^{6}$. Right: The black line in the $C_{H\Box}$ vs $C_{H}$ plane corresponds to the CQ model.
@@
where $\lambda_{\rm SM}$ is the SM Higgs quartic. The operator $\partial_{\mu}|H|^{2}\partial^{\mu}|H|^{2}$ is equivalent to $Q_{H\Box}$ via integration by parts. Furthermore, the terms containing $D^{2}H$ can be reduced using the Higgs equation of motion to another $|H|^{6}$ term with coefficient $2\lambda_{\rm SM}\lambda_{4}^{2}/(3\pi^{2}M_{4}^{2})$ plus dimension-6 Yukawa corrections of the form $\sum_{\psi}Y_{\rm\psi}^{ij}|H|^{2}\bar{\psi}_{L}^{i}H\psi_{R}^{j}$. These terms correct Higgs decays, but we drop them in what follows since they are always suppressed by small SM Yukawa couplings $Y_{\rm\psi}$. Similarly, the total loop correction to $|H|^{6}$ proportional to $\lambda_{\rm SM}$ is of size $10^{-2}$ relative to the tree contribution and can be safely neglected. Thus for phenomenological purposes we may work in the limit where we keep matching contributions only in $\lambda_{4},g_{1}$, and $g_{2}$. We note that the $\lambda_{4}$ terms of our result agree with those of Durieux:2022hbu , while the contributions to $\mathcal{O}_{2B,2W,3W}$ can be easily derived from Eq. 35.
In Fig. 13 (left), we show the sensitivity of FCC-ee to the custodial weak quadruplet model on and off the $Z/W$ pole, where we run from a fixed high-scale $\Lambda=$1\text{\,}\mathrm{TeV}$\$ for simplicity. We compare this with the above-pole sensitivity obtained only by considering the tree-level modification of the Higgs self-coupling (dashed blue line). We see that the full above-pole result shows a significant reduction in sensitivity. This is because while the contribution to $\mathcal{C}_{H\Box}$ is loop suppressed and $\mathcal{C}_{H}$ is generated at tree level, $\mathcal{C}_{H\Box}$ enters $\sigma(ZH)$ at LO while $\mathcal{C}_{H}$ enters at NLO, with a relative minus sign. The CQ model predicts a fixed ratio of $\mathcal{C}_{H}/\mathcal{C}_{H\Box}=+8\pi^{2}/3$, so the two contributions partially cancel in the $ZH$ cross section. This can be seen in Fig. 13 (right), where the CQ model accidentally approximately aligns with the flattest above-pole direction. On the other hand, the $Z$-pole provides a strong bound in this direction, dominantly via the NLO contribution of $\mathcal{C}_{H\Box}$ to $\mathcal{C}_{HD}$. Finally, the $\lambda_{4}$-independent lower bound of $M_{4}\gtrsim 800$ GeV comes from the $\hat{W}$ and $\hat{Y}$ parameters, which scale as $g_{\rm EW}^{2}/M_{4}^{2}$.
@@
For $\lambda_{4}=1$ and $M_{4}=1$ TeV, this yields a contribution to the $T$ parameter of $\hat{T}=-v^{2}\mathcal{C}_{HD}(m_{Z})/2\approx 10^{-5}$, which is at the detectable level according to Eq. 23. However, $Q_{HD}$ is not the only operator activated (see e.g. Eq. 11), and using Eq. 34 we find $M_{4}/\lambda_{4}\gtrsim 1.5$ TeV, in agreement with the on-pole limit in Fig. 13.
@@ 5 Discussion and conclusions
We have emphasised in this work how high-precision accuracy at Tera-$Z$ complements and can even beat high-energy measurements. When they are sensitive to the same new physics modifications entering at different orders or with energy-growing effects, the increased statistics compensate for higher loop suppression on the $Z$ pole or for energy enhancements above-pole. It is a simple counting exercise to see that $Z$-pole sensitivity at NLO is of the same order of magnitude as above-pole sensitivity at LO. Tera-$Z$ is therefore expected to provide competitive NLO sensitivity to above-pole measurements at LO for any Wilson coefficient that they both have access to. This “accuracy complements energy” principle also applies at NNLO on-pole vs. NLO above-pole, as well as if a Wilson coefficient enters at LO both on-pole and above-pole but is energy-enhanced in the latter.
@@
We studied new physics contributions to the Higgs 2-point and 3-point functions, which modify the Higgs self-energy, self-coupling, and Higgs-vector-vector interactions, and to the gauge boson 2-point and 3-point functions which enter in the oblique parameters and anomalous triple-gauge boson couplings. The Higgs self-energy and vector-vector interactions contribute at leading order to $e^{+}e^{-}\to ZH$, while the Higgs self-coupling enters this process at NLO. On the other hand, they all affect the $Z$ pole oblique parameters at NLO and NNLO respectively, leading as expected to competitive on-pole projected sensitivities at FCC-ee and complementary bounds when combined with above-pole data. We find a similar story for the anomalous triple-gauge couplings that enter at leading order in $WW$ and at NLO at the $Z$ pole. Finally, the $Z$ pole oblique parameters $\hat{W}$ and $\hat{Y}$ can be constrained at leading order both on- and above-pole, with the former being statistics-enhanced and the latter energy-enhanced, such that they also lead to an overall similar and complementary projected bounds.
@@
Since so many Wilson coefficients enter the oblique parameters at higher orders, global SMEFT fits beyond leading order will find large correlations between seemingly unrelated above-pole Wilson coefficients, and cancellations between them when marginalising will give the appearance of weaker bounds. Hence it is crucial to emphasise that with Tera-$Z$ it will be especially important to match to explicit UV models to extract useful information, as the dynamical information from the model will correlate operators and pick out specific directions in the SMEFT Wilson coefficient space that typically result in stronger bounds instead. The fundamental reason is because the volume of the region that can be constrained in Wilson coefficient space is reduced by combining datasets beyond leading order. We find that this is indeed the case for the $Z_{2}$-symmetric real singlet scalar, scalar and fermion WIMPs, and a custodial weak quadruplet scalar. These explicit UV models all benefit significantly from the inclusion of on-pole data expanding coverage of their projected parameter space at FCC-ee.
@@ Appendix A SMEFT theory expression for $\sigma(e^{+}e^{-}\rightarrow W^{+}W^{-}$)
In this appendix, we give the results of our calculation of $\sigma(e^{+}e^{-}\rightarrow W^{+}W^{-})$ at tree level for the relevant center of mass energies for FCC-ee, $E_{\rm{CM}}=$240\text{\,}\mathrm{GeV}$\$ and $365\text{\,}\mathrm{GeV}$. We calculate this cross section using the Feynman Rules implemented in SMEFT-FR Dedes:2017zog ; Dedes:2019uzs ; Dedes:2023zws , from which we generate a FeynArts Hahn:2000kx model, that can then be imported into FeynCalc Mertig:1990an ; Shtabovenko:2016sxi ; Shtabovenko:2020gxv ; Shtabovenko:2023idz to calculate the final cross section, using the $\{M_{W},M_{Z},G_{F}\}$-scheme. We present our results first in the non-redundant Warsaw Basis defined in the WCXF-format Aebischer:2017ugx and make no flavor assumptions on the coefficients. With these choices, only 8 operators modify the cross section:
@@
$E_{\rm{CM}}$ [GeV] SM $\mathrm{pb}$] $\mathcal{C}_{W}$ $\mathcal{C}_{HWB}$ $\mathcal{C}_{HD}$ $\mathcal{C}_{He}]_{11} [$$\mathcal{C}_{Hl}^{(1)}]_{11} [$$\mathcal{C}_{Hl}^{(3)}]_{11} [$$\mathcal{C}_{Hl}^{(3)}]_{22} [$$\mathcal{C}_{ll}]_{1221}$
@@
Table 2: Contribution $b_{i}\cdot 10^{3}$ of Wilson Coefficients $C_{i}/($1\text{\,}\mathrm{TeV}$)^{2}$ to $\sigma(e^{+}e^{-}\rightarrow W^{+}W^{-})$ normalised to the Standard Model value given in the second row for $E_{CM}=$240\text{\,}\text{GeV}$\$ and $365\text{\,}\text{GeV}$, corresponding to the $ZH$- and $t\bar{t}$-thresholds.
As described in Sec. 3.2, it is sometimes more convenient to work in the SILH basis. Both share a large number of coefficients, but motivated by Higgs and gauge coupling modifications, as well as the situation described in Stefanek:2024kds , we focus on the subset of coefficients introduced in Eqs. 17-20 and 25 -27 and App. B.2. With these assumptions, the only operators affecting the diboson production cross section:
@@
Their relative contribution assuming $\Lambda=$1\text{\,}\mathrm{TeV}$\$ is given in Table 3.
$E_{\rm{CM}}$ [GeV] SM $\mathrm{pb}$] $\mathcal{S}_{3W}$ $\mathcal{S}_{HB}$ $\mathcal{S}_{HW}$ $\mathcal{S}_{2B}$ $\mathcal{S}_{2W}$ $\mathcal{S}_{B+W}$ $\mathcal{S}_{B-W}$ $\mathcal{S}_{T}$
@@ Appendix B Single operator sensitivities
We report here the projected sensitivity of FCC-ee to SMEFT Wilson coeffficients (WCs) renormalised at $\Lambda=$1\text{\,}\mathrm{TeV}$\$ when activating one WC at a time using the likelihood described in 2. Since we are setting the projected experimental values to their SM value and our theory predictions are linear in $\mathcal{C}_{i}$, our likelihood is Gaussian and centered at $\mathcal{C}_{i}=0$. Thus, we quote the $2\sigma$ intervals in terms of $\Lambda_{\rm{eff}}^{\rm{i}}=\left|\left(\sqrt{\mathcal{C}_{i}}\right)^{-1}\right|$ using observables measured on the $Z$ and $W$ pole (on-pole), as well as above-pole observables measured at the $ZH$ and $t\bar{t}$ thresholds (above-pole) and, lastly, their combined sensitivity. For the on- and above-pole observables, we quote also the observables with the largest pull
@@
with $\mathcal{C}_{j}$ evaluated at the $2\sigma$ limit. This can be interpreted as the observable that most strongly constrains the coefficient. For the combined sensitivity, since on- and above-pole are not correlated, the strongest bound gives the largest pull as well.